OpenStudy (anonymous):

Consider the function y = 3x5 – 25x3 + 60x + 1. Use the second derivative test to determine whether our function has a relative maximum or minimum at x = –1. Which of the following describes what you found? (Points : 1) x = –1 isn't a critical point, so I can't use the second derivative test, and there's neither a max nor a min there. there, so there is a relative minimum at x = –1. there, so there is a relative maximum at x = –1. there, so there is neither a max nor a min there. there, so the second derivative test fails, and I'd have to use a

OpenStudy (campbell_st):

why not just reply in the question you posted... rather than reposting 3 times...

OpenStudy (anonymous):

ITS A DIFFERENT QUESTION

OpenStudy (campbell_st):

wow...same curve... and given I gave you the factored form... who would have known...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

well you have a slight error given the posting you just closed f′(x)=15x4−75x2+60 which can be factored to f′(x)=15(x4−5x2+4) which becomes f′(x)(x2−1)(x2−4) or f′(x)=(x−1)(x+1)(x−2)(x+2) now you can play witht he 2nd derivative to check their nature

OpenStudy (anonymous):

THANKYOU

OpenStudy (anonymous):

MY BAD

OpenStudy (anonymous):

Consider the function y = 3x5 – 25x3 + 60x + 1. Use the first derivative test to decide whether this function has a maximum at x = 1. Which of the following describes what you found? (Points : 1) The derivative is positive to the left of x = 1 and negative to the right of x = 1, so the function has a relative minimum at x = 1. The derivative is positive to the left of x = 1 and negative to the right of x = 1, so the function has a relative maximum at x = 1. The derivative is positive to the left of x = 1 and positive to the right of x = 1, so the function has neither a relative maximum nor a minimum at x = 1. The derivative is negative to the left of x = 1 and positive to the right of x = 1, so the function has a relative minimum at x = 1. None of these apply.

OpenStudy (campbell_st):

here is a graph of your curve with with all the stationary points and points of inflexion shown.... oh... and infexion can also be inflection... hope this helps with the new question

OpenStudy (anonymous):

ha! im working the same thing. my technique was using a number lick. pick a point aside of the critical point and see if the function is positive or negative, then you'll be able to tell if it's a max or min.